# Global regularity for solutions of magnetohydrodynamic equations with large initial data

### v7

@@ 1. Introduction
(2.58) $\frac{f(r)}{r-2j}$ $=$ $\frac{h(r)f(r)}{h(r)(r-2j)}$
$=$ $\frac{1}{h(r)(r-2j)}\left[h(q)f(q)+\int_{q}^{r}\left(h(s)f(s)\right)^{\prime}ds\right]$
$=$ $\frac{1}{h(r)(r-2j)}\left[\int_{q}^{r}h(s)f^{\prime}(s)ds+\int_{q}^{r}h^{\prime}(s)f(s)ds\right].$
To estimate $f^{\prime}(s)$, we appeal to (2.34). But we will use a slightly different version of it. Without any loss of generality, we may assume that the second term in (2.34) is less that or equal to the third term there, i.e.,
$16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}\leq c_{3}\|w\|_{L_{N},Q_{T}}^{s_{0}}\|w\|_{s,Q_{T}}^{\frac{s}{s-M_{q}}}\ \ \mbox{for each $s>M_{q}$}.$
(We have changed $r$ in (2.34) to $s$.) Were this not true, we would have nothing more to prove. Consequently, (2.34) is reduced to
$\|w\|_{\infty,Q_{T}}$ $\leq$ $2c_{3}\|w\|_{L_{N},Q_{T}}^{s_{0}}\|w\|_{s,Q_{T}}^{\frac{s}{s-M_{q}}}\ \ \mbox{for $s\in(M_{q},\infty)$}.$
This together with (2.45) implies that
$f^{\prime}(s)\leq\ln\left(\|w\|_{\infty,Q_{T}}\right)\leq\ln\left(2c_{3}\|w\|_{L_{N},Q_{T}}^{s_{0}}\right)+\frac{f(s)}{s-M_{q}},\ \ s>M_{q}.$
In view of (2.49), (2.55), and (2.53), we have
(2.59) $\int_{q}^{r}h(s)f^{\prime}(s)ds+\int_{q}^{r}h^{\prime}(s)f(s)ds$
$\leq$ $\ln\left(2c_{3}\|w\|_{L_{N},Q_{T}}^{s_{0}}\right)\int_{q}^{r}h(s)ds+\int_{q}^{r}\left(\frac{h(s)}{s-M_{q}}+h^{\prime}(s)\right)f(s)ds$
$=$ $\ln\left(2c_{3}\|w\|_{L_{N},Q_{T}}^{s_{0}}\right)\int_{q}^{r}h(s)ds+\int_{q}^{\xi}f(s)ds-A\int_{\xi}^{r}f(s)ds.$
Fix
(2.60) $d\in(M_{q},q).$
Consider the line through $(d,f(d))$ and $(\xi,f(\xi))$. The convexity of $f(s)$ implies
$f(s)$ $\leq$ $f(\xi)+\frac{f(\xi)-f(d)}{\xi-d}(s-\xi)\ \ \mbox{for $s\in(q,\xi)$},$
$f(s)$ $\geq$ $f(\xi)+\frac{f(\xi)-f(d)}{\xi-d}(s-\xi)\ \ \mbox{for $s\in(\xi,r)$.}$
Consequently,
(2.61) $\int_{q}^{\xi}f(s)ds-A\int_{\xi}^{r}f(s)ds$
$\leq$ $f(\xi)(\xi-q)-\frac{(f(\xi)-f(d))(\xi-q)^{2}}{2(\xi-d)}-Af(\xi)(r-\xi)-\frac{A(f(\xi)-f(d))(r-\xi)^{2}}{2(\xi-d)}$
$=$ $f(\xi)\left[\xi-q-A(r-\xi)-\frac{(\xi-q)^{2}+A(r-\xi)^{2}}{2(\xi-d)}\right]+\frac{f(d)\left[(\xi-q)^{2}+A(r-\xi)^{2}\right]}{2(\xi-d)}.$
We pick $A$ so that
(2.62) $\xi-q-A(r-\xi)-\frac{(\xi-q)^{2}+A(r-\xi)^{2}}{2(\xi-d)}=0.$
This yields
(2.63) $A=\frac{\xi-q-\frac{(\xi-q)^{2}}{2(\xi-d)}}{r-\xi+\frac{(r-\xi)^{2}}{2(\xi-d)}}=\frac{(\xi-q)(\xi+q-2d)}{(r-\xi)(r+\xi-2d)}.$
We easily see from (2.60) that $A$ so chosen satisfies (2.57), which, in turn, results in (2.55). We can also confirm this by directly substituting (2.63) into the expression for $h(r)$ in (2.56), thereby obtaining
(2.64) $h(r)$ $=$ $\frac{(\xi-q)\left[-(\xi+q-2d)(r+\xi-2M_{q})+(r+\xi-2d)(\xi+q-2M_{q})\right]}{2(r-M_{q})(r+\xi-2d)}$
$=$ $\frac{(\xi-q)[(d-M_{q})r+M_{q}q-d^{2}]}{(r-M_{q})(r+\xi-2d)}>0.$
The last step here is due to (2.60). Now we incorporate (2.62) and (2.63) into (2.61) to deduce
$\int_{q}^{\xi}f(s)ds-A\int_{\xi}^{r}f(s)ds$ $\leq$ $f(d)\left[\xi-q-A(r-\xi)\right]$
$=$ $\frac{f(d)(\xi-q)(r-q)}{r+\xi-2d}.$
Collecting this in (2.59), we arrive at
$\int_{q}^{r}h(s)f^{\prime}(s)ds+\int_{q}^{r}h^{\prime}(s)f(s)ds$
$\leq$ $\ln\left(2c_{3}\|w\|_{L_{N},Q_{T}}^{s_{0}}\right)\int_{q}^{r}h(s)ds+\frac{f(d)(\xi-q)(r-q)}{r+\xi-2d}.$
Use this in (2.58) to obtain
(2.65) $\frac{f(r)}{r-2j}$ $\leq$ $\frac{1}{h(r)(r-2j)}\left[\ln\left(2c_{3}\|w\|_{L_{N},Q_{T}}^{s_{0}}\right)\int_{q}^{r}h(s)ds+\frac{f(d)(\xi-q)(r-q)}{r+\xi-2d}\right].$
On the other hand, we can infer from (2.46) that
$-\frac{jf^{\prime}(r)}{r-2j}\leq-\frac{jf^{\prime}(q)}{r-2j}\leq-\frac{j(f(q)-f(L_{N}))}{(r-2j)(q-L_{N})}.$
Collect this and (2.65) in (2.48) obtain
(2.66) $\|w\|_{\infty,Q_{T}}\leq 16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}+c\|w\|_{L_{N},Q_{T}}^{s_{3}}\|w\|_{q,Q_{T}}^{H(r)}\|w\|_{d,Q_{T}}^{\frac{d(\xi-q)(r-q)}{h(r)(r-2j)(r+\xi-2d)}},$
where
(2.67) $H(r)$ $=$ $\frac{q(M_{q}-j)}{(q-L_{N})(r-2j)}-\frac{jq}{(q-L_{N})\left(r-2j\right)},$
$s_{3}$ $=$ $s_{1}+\frac{jL_{N}}{(q-L_{N})\left(r-2j\right)}+\frac{s_{0}\int_{q}^{r}h(s)ds}{h(r)(r-2j)}.$
In view of (2.60), we can form the interpolation inequality
$\|w\|_{d,Q_{T}}\leq\|w\|_{q,Q_{T}}^{\frac{q(d-L_{N})}{d(q-L_{N})}}\|w\|_{L_{N},Q_{T}}^{\frac{L_{N}(q-d)}{d(q-L_{N})}}.$
Incorporate this into (2.66) to obtain
(2.68) $\|w\|_{\infty,Q_{T}}\leq 16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}+c\|w\|_{L_{N},Q_{T}}^{s_{3}+\frac{(\xi-q)(r-q)L_{N}(q-d)}{h(r)(r-2j)(r+\xi-2d)(q-L_{N})}}\|w\|_{q,Q_{T}}^{H(r)+\frac{(\xi-q)(r-q)q(d-L_{N})}{h(r)(r-2j)(r+\xi-2d)(q-L_{N})}}.$
We must verify that
(2.69) $H(r)+\frac{(\xi-q)(r-q)q(d-L_{N})}{h(r)(r-2j)(r+\xi-2d)(q-L_{N})}<\frac{q}{q-L_{N}}.$
Plug (2.67) and (2.64) into this to derive
$M_{q}+\frac{(r-q)(d-L_{N})(r-M_{q})}{(d-M_{q})r+M_{q}q-d^{2}}<r.$
In view of (2.43) and (2.60), we can solve this inequality for $r$ to obtain
(2.70) $r<\frac{(d+M_{q}-L_{N})q-d^{2}}{M_{q}-L_{N}}.$
This condition must be consistent with (2.50). That is, there must hold
(2.71) $\frac{(d+M_{q}-L_{N})q-d^{2}}{M_{q}-L_{N}}>q.$
This inequality is an easy consequence of (2.60).
In summary, the order in which we pick our parameters is as follows: We first fix $q$ so that both (2.16) and (2.49) hold. Then pick $d$ as in (2.60), from which (2.71) follows. As a result, we can take
$r\in\left(q,\frac{(d+M_{q}-L_{N})q-d^{2}}{M_{q}-L_{N}}\right).$
This implies both (2.70) and (2.50). Now we can define $A$ as in (2.63) for $\xi\in(q,r)$. Finally, $j$ can be any number in the interval $\left(0,\frac{q}{2}\right)$. Hence, (2.22) is valid.
@@
(2.72) $\|w\|_{q,Q_{T}}>1\ \ \mbox{for each $q>M_{q}$.}$
Otherwise, (2.34) would be enough to imply our theorem. Under (2.72), we may suppose
(2.73) $H(r)+\frac{(\xi-q)(r-q)q(d-L_{N})}{h(r)(r-2j)(r+\xi-2d)(q-L_{N})}>0.$
Were this not true, (2.72) combined with (2.68) would yield our theorem. We can form the interpolation inequality
$\|w\|_{q,Q_{T}}\leq\|w\|_{\infty,Q_{T}}^{\frac{q-L_{N}}{q}}\|w\|_{L_{N},Q_{T}}^{\frac{L_{N}}{q}}.$
Collect this in (2.68) and keep (2.73) in mind to get
$\|w\|_{\infty,Q_{T}}$ $\leq$ $16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}+c\|w\|_{L_{N},Q_{T}}^{s_{4}}\|w\|_{\infty,Q_{T}}^{\theta},$
where
$\theta$ $=$ $\frac{(q-L_{N})}{q}\left[H(r)+\frac{(\xi-q)(r-q)q(d-L_{N})}{h(r)(r-2j)(r+\xi-2d)(q-L_{N})}\right],$
$s_{4}$ $=$ $s_{3}+\frac{(\xi-q)(r-q)L_{N}(q-d)}{h(r)(r-2j)(r+\xi-2d)(q-L_{N})}$
$+\frac{L_{N}}{q}\left[H(r)+\frac{(\xi-q)(r-q)q(d-L_{N})}{h(r)(r-2j)(r+\xi-2d)(q-L_{N})}\right].$
According to (2.69) and (2.73), we have
$\theta\in(0,1).$

### v8

@@ 1. Introduction
$\frac{r-2j+s_{j}}{2j+\varepsilon-L_{N}}$ $=$ $\frac{r-j-\frac{(2j+\varepsilon-q)(j-M_{q})}{q-L_{N}}}{2j+\varepsilon-L_{N}}$
$=$ $\frac{r-M_{q}-\frac{(2j+\varepsilon-L_{N})(j-M_{q})}{q-L_{N}}}{2j+\varepsilon-L_{N}}$
$=$ $\frac{r-M_{q}}{2j+\varepsilon-L_{N}}-\frac{(j-M_{q})}{q-L_{N}}.$
By the same token,
$\frac{2j-s_{j}-M_{q}}{2j+\varepsilon-L_{N}}$ $=$ $\frac{j+\frac{(2j+\varepsilon-q)(j-M_{q})}{q-L_{N}}-M_{q}}{2j+\varepsilon-L_{N}}$
$=$ $\frac{(j-M_{q})}{q-L_{N}},$
$\frac{s_{j}+\varepsilon}{2j+\varepsilon-L_{N}}$ $=$ $\frac{j-\frac{(2j+\varepsilon-q)(j-M_{q})}{q-L_{N}}+\varepsilon}{2j+\varepsilon-L_{N}}$
$=$ $\frac{2j-M_{q}-\frac{(2j+\varepsilon-L_{N})(j-M_{q})}{q-L_{N}}+\varepsilon}{2j+\varepsilon-L_{N}}$
$=$ $\frac{2j-M_{q}+\varepsilon}{2j+\varepsilon-L_{N}}-\frac{j-M_{q}}{q-L_{N}}.$
Incorporate the preceding three equations into the expression for $h(r)$ in (2.70) to derive
$h(r)$ $=$ $\frac{(r-2j)(r-L_{N})}{2j+\varepsilon-L_{N}}+\frac{(j-M_{q})(r-L_{N})}{(q-L_{N})}-r+M_{q}$
$+(r-2j-\varepsilon-s_{j})\left[e^{\frac{r-2j-\varepsilon}{s_{j}}}\left[\frac{2j-M_{q}+\varepsilon}{2j+\varepsilon-L_{N}}-1\right]-\frac{r-M_{q}}{2j+\varepsilon-L_{N}}\right]$
$=$ $\frac{(r-2j)(r-L_{N})}{2j+\varepsilon-L_{N}}+\frac{(j-M_{q})(r-L_{N})}{(q-L_{N})}-\frac{(r-M_{q})(r-s_{j}-L_{N})}{2j+\varepsilon-L_{N}}$
$-\frac{(r-2j-\varepsilon-s_{j})(M_{q}-L_{N})e^{\frac{r-2j-\varepsilon}{s_{j}}}}{2j+\varepsilon-L_{N}}$
$=$ $\frac{-(2j-s_{j}-M_{q})r+L_{N}(2j-M_{q})}{2j+\varepsilon-L_{N}}+\frac{(j-M_{q})(r-L_{N})}{(q-L_{N})}$
$-\frac{(r-2j-\varepsilon-s_{j})(M_{q}-L_{N})e^{\frac{r-2j-\varepsilon}{s_{j}}}}{2j+\varepsilon-L_{N}}$
$=$ $\frac{L_{N}(2j-M_{q})}{2j+\varepsilon-L_{N}}-\frac{L_{N}(j-M_{q})}{(q-L_{N})}-\frac{(r-2j-\varepsilon-s_{j})(M_{q}-L_{N})e^{\frac{r-2j-\varepsilon}{s_{j}}}}{2j+\varepsilon-L_{N}}.$
In view of (2.42), for (2.71) to hold, we must have
$\frac{(2j-M_{q})}{2j+\varepsilon-L_{N}}-\frac{(j-M_{q})}{(q-L_{N})}>0.$
Solve this for $\varepsilon$ to deduce
$\varepsilon$ $<$ $\frac{q-L_{N}}{j-M_{q}}\left[2j-M_{q}-\frac{(j-M_{q})(2j-L_{N})}{q-L_{N}}\right]$
$=$ $\frac{q-L_{N}}{j-M_{q}}\left[j-\frac{(j-M_{q})(2j-q)}{q-L_{N}}\right],$
which is exactly our assumption (2.61).
In summary, the order in which we choose our parameters is as follows: Let $q$ be given as in (2.50). By virtue of (2.63), we may take
$j\in\left(\max\left\{\frac{q}{2},j_{1}\right\},j_{2}\right).$
This implies (2.51) and enables us to select $\varepsilon$ as in (2.61), which, in turn, guarantees (2.71) for $r$ close to $2j+\varepsilon+s_{j}$ from the right-hand side. That is, we have both (2.66) and (2.71). Equipped with this, we can choose $\eta\in(0,1)$ so that (2.70) is satisfied.
@@
(2.72) $\|w\|_{2j+\varepsilon,Q_{T}}>1.$
Otherwise, (2.34) would be enough to imply our theorem. Under (2.72), we may suppose that the last exponent $\beta_{2}$ in (2.67) is positive. Were this not true, (2.72) combined with (2.67) would yield our theorem. In view of (2.16) and (2.41), we can form the interpolation inequality
$\|w\|_{2j+\varepsilon,Q_{T}}\leq\|w\|_{\infty,Q_{T}}^{\frac{2j+\varepsilon-L_{N}}{2j+\varepsilon}}\|w\|_{L_{N},Q_{T}}^{\frac{L_{N}}{2j+\varepsilon}}.$
Collect this in (2.67) and keep in mind that $\beta_{2}>0$ to get
$\|w\|_{\infty,Q_{T}}$ $\leq$ $16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}+c\|w\|_{L_{N},Q_{T}}^{s_{1}+\frac{L_{N}\beta_{2}}{2j+\varepsilon}}\|w\|_{\infty,Q_{T}}^{\frac{(2j+\varepsilon-L_{N})\beta_{2}}{2j+\varepsilon}},$
According to (2.69), we have
$\frac{(2j+\varepsilon-L_{N})\beta_{2}}{2j+\varepsilon}\in(0,1).$